OpenStudy (jmartinez638):

Calculate the following series:

OpenStudy (jmartinez638):

\[\sum_{k=1}^{20} 3k^2 +5k\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

break it apart first

OpenStudy (misty1212):

\[3\sum k^2+5\sum k\] then use the formulas for each

OpenStudy (jmartinez638):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n} (4+2k/3\times1/n)\times1/n\]

OpenStudy (jmartinez638):

Hiya! ;)

OpenStudy (misty1212):

i can't read the second one

OpenStudy (jmartinez638):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n} (1+2k/n)^8 \times1/n\]

OpenStudy (jmartinez638):

One sec, I will edit it.

OpenStudy (misty1212):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n} (4+\frac{2k}{3})\times\frac{1}{n})\times\frac{1}{n}\] is my guess

OpenStudy (misty1212):

some sorta integral right?

OpenStudy (misty1212):

turn it to \[\frac{1}{n^2}\sum 4+\frac{2}{3}\sum k\]

OpenStudy (misty1212):

actually \[\frac{1}{n^2}\left(\sum_{k=1}^n 4+\frac{2}{3}\sum_{k=1}^n k\right)\]

OpenStudy (misty1212):

unless am reading it wrong, which is possible

OpenStudy (zarkon):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n} (1+2k/n)^8 \times1/n\] are you supposed to write this as an integral and evaluate?

OpenStudy (jmartinez638):

Just says evaluate, your thought is as good as mine.

OpenStudy (jmartinez638):

Not to be blunt, I just don't know